If a line contains the point (0, -1) and has a slope of 2, then which of the following points also lies on the line?
(2, 1)
(1, 1)
(0, 1)

Answer by richwmiller(9135)

y=mx+b
with a slope of 2 we now have
y=2x+b
since the point is (0,-1)
we plug them in and solve for b
-1=2(0)+b
-1=b
so now we have y=2x-1
so now we just plug in the ordered pairs to see which give equalities
(2, 1)
y=2x-1
1=2*2-1
1=4-1
1=3
not on the line
(1, 1)
y=2x-1
1=2*1-1
1=2-1
1=1 yes on the line
(0, 1)
y=2x-1
1=2*0-1
1=-1
not on the line